# Burnside type results for fusion categories

### v10

@@ 10.1. Modular category case
In this section let $\mathcal{C}$ be a weakly-integral modular fusion category. We apply the results of previous section to $K(\mathcal{C})$ and prove Theorem 1.13 and Theorem 1.16. Recall the definition of $\mathcal{J}_{{}_{\mathcal{D}}}$ from Equation (9.7).
Composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the natural isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4) one obtains an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$. This can also be written as an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$. It was shown in [Bur21, Theorem 4.1] that in fact this is an isomorphism of normalized hypergroup ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}.$ Recall that $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$.
@@
This isomorphism establishes a canonical bijection between the sets $\mathcal{I}$ and $\mathcal{J}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$ for all $i\in\mathcal{I}$.
Through the rest of the paper we identify $\mathcal{I}$ and $\mathcal{J}$ via the above bijection.
@@
As mentioned above ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$. Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if ${\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
@@
(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$
@@
The result follows since $({\mathcal{C}_{\mathrm{ad}}})^{\prime}={\mathcal{C}_{\mathrm{pt}}}$. Then $X_{i}\in\mathrm{Irr}(({\mathcal{C}_{\mathrm{ad}}})^{\prime})$ if and only if $i\in J_{\mathcal{C}_{\mathrm{ad}}}$ by the second item of [Bur20, Theorem 6.2]. ∎
@@
(10.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(x_{j})F_{j}$
@@
Since $K(\mathcal{C})$ is Burnside we may apply ${\widehat{\mathrm{f}_{Q}}}^{-1}$ to Equation (9.3). Under the above bijection between $\mathcal{I}$ and $\mathcal{J}$ one has ${\widehat{\mathrm{f}_{Q}}}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, see [Bur20, Theorem 6.1]. Thus the above Equation (9.3) becomes
$\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
@@
Squaring Equation (10.8) one obtains that $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$. Note that $\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$ by the definition of ${J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$.∎
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside.
@@
Let $\mathcal{C}$ be a fusion category such that the dual ${\widehat{{K(\mathcal{C})}}}$ is a Burnside hypergroup. Then
$\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
@@
For any category $\mathcal{C}$ define $P_{\mathcal{C}}:=\prod_{i=1}^{m}d_{i}$. Suppose that $\big{(}\prod_{i=1}^{m}x_{i}\big{)}^{2}=\sum_{i=1}^{m}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$. Then equalizing coefficients in Equation (1.7) one obtains the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. For $i=1$ it follows that $n_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility. ∎
@@
By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral. Note that since ${\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6. Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})$. ∎
@@
$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
@@
The first divisibility follows from Proposition (10.6). Suppose now that $\mathcal{C}$ is a weakly-integral modular fusion category. By [ENO05, Proposition 8.27] one has that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer. By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$. The second item follows since for a modular category one has $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$. On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})\big{)}$.
Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$ and $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $1\leq i\leq m$. ∎
Proof of Theorem 1.17
Proof.
Since $g.c.d(d_{i}^{2},\;d)=1$ it follows from Equation (1.14) that $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Let now ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ a pointed fusion subcategory such that $\mathrm{FPdim}({\mathcal{D}})=d$. Since $\mathrm{FPdim}({\mathcal{D}}^{\prime})=N$ and $\mathrm{g.c.d}(d,N)=1$ it follows that ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$ and therefore ${\mathcal{D}}$ is non-degenerate. Thus $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$ and ${\mathcal{D}}^{\prime}$ is non-degenerate as well. ∎
@@
Note that the above Corollary improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
@@
Immediate from Theorem 1.17, we just need to check that $g.c.d(d_{i}^{2},\;d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a square-free factor. ∎
Lemma 10.9.
With the above notations, for an integral perfect modular category, $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ and
$d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})\mid P_{\mathcal{C}}^{2}$
Proof.
Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, see Remark 11.13. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ by Theorem 1.16. Since $d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})$ the result follows. ∎
Lemma 10.9 extends trivially to the weakly-integral case because:
@@
By the proof of Lemma 10.9 then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$. But [EGNO15, Exercice 9.6.12] states that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
@@
Let $\mathcal{C}$ be an integral perfect modular fusion category. If $p$ is a prime number dividing $\mathrm{FPdim}(\mathcal{C})$ then $p^{2}\mid\mathrm{FPdim}(\mathcal{C})$.
Corollary 10.12.
Any integral modular category $\mathcal{C}$ such that there is a prime $p$ whose exponent in $\mathrm{FPdim}(\mathcal{C})$ is $1$ (i.e $m\neq 1$) it is not perfect.
Corollary 10.13.
If $\mathcal{C}$ is a perfect, integral, even dimensional modular fusion category then $4\mid\mathrm{FPdim}(\mathcal{C})$.
10.2. On the Statements (1.21) and (1.22) and Conjecture (1.20)
In [LPR23, Theorem 5.6] it was shown that the statements (1.21) and (1.22) are equivalent. Next we will show that any of these statements implies Conjecture 1.20 formulated in [CP22].
@@
Existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
@@
Suppose that there exists a perfect modular odd-dimensional fusion category $\mathcal{C}$.
Note first that $\mathcal{C}$ has no pointed subcategories since it is perfect. It also has no symmetric subcategories, since then these subcategories are of the type $\mathrm{Rep}(G)$ with $G$ a perfect group. But then $|G|$ is divisible by $4$ and this is a contradiction since the $\mathrm{FPdim}(\mathcal{C})$ is odd.
On the other hand, by a dimension argument, $\mathcal{C}$ has at least one minimal subcategory under the inclusion (and hence a simple subcategory), say ${\mathcal{D}}$. If ${\mathcal{D}}$ is a modular fusion category then ${\mathcal{D}}$ is an example of a non-pointed simple integral modular fusion category.
If ${\mathcal{D}}$ is not modular then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is a non-trivial symmetric fusion subcategory which as explained above, it is not possible. ∎

### v11

@@ 10.1. Modular category case
In this section, let $\mathcal{C}$ denote a modular fusion category. We will utilize the results from the previous section to $K(\mathcal{C})$ and prove Theorems 1.16 and 1.19. Recall the definition of $\mathcal{I}_{{}_{\mathcal{D}}}$ as provided in Definition (6.3).
By composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4), we obtain an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$. This can also be expressed as an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$. It was shown in [Bur21, Remark 4.2] that this is, in fact, an isomorphism of normalized hypergroups: $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$, so that $\widehat{\mathrm{f}}_{Q}(\frac{x_{i}}{d_{i}})=\mu_{i}$.
@@
Drinfeld’s map $\mathrm{f}_{Q}$ is an algebra isomorphism since the S-matrix is invertible. It establishes a canonical bijection between the index sets of $\{F_{i}\}$ and $\{E_{i}\}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$, for all $i\in\mathcal{I}$.
@@
As mentioned above $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup. Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if $^{\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
@@
(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}$
@@
Apply Theorem 6.5 with the isomorphism of normalized hypergroups $\widehat{\mathrm{f}}_{Q}$. ∎
@@
(10.8) $\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}{\mathrm{sgn}}(x_{j})F_{j}$
@@
Since $K(\mathcal{C})$ is Burnside, we can apply $\widehat{\mathrm{f}}_{Q}^{-1}$ to Equation (9.3), but $\widehat{\mathrm{f}}_{Q}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, so we get
$\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
@@
Squaring Equation (10.8) we obtain that $\big{(}\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}$ by the definition of ${{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$. The result follows by Equation (5.2).∎
Recall that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
@@
Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and the dual $^{\widehat{{K(\mathcal{C})}}}$ is Burnside. Then
$\frac{(\prod_{i\in\mathcal{I}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
@@
Let $P_{\mathcal{C}}$ be $\prod_{i\in\mathcal{I}}d_{i}$. Suppose that $\big{(}\prod_{i\in\mathcal{I}}x_{i}\big{)}^{2}=\sum_{i\in\mathcal{I}}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$. Then equalizing coefficients in Equation (1.6), we obtain the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Take $i=1$ then $N_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility. ∎
@@
By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral. Note that since $^{\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6. Thus ${\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$ is an integer (as a rational algebraic integer). Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})$. ∎
@@
$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}).$
Recall that a weakly-integral fusion category with a commutative Grothendieck ring is Burnside by [Bur23, Theorem 2]. Moreover, a Burnside modular fusion category is dual-Burnside by Corollary 10.3. So a weakly-integral modular fusion category is dual-Burnside. Proof of Theorem 1.19
Proof.
The first divisibility follows from Proposition (10.6). Now, let $\mathcal{C}$ be a weakly-integral modular fusion category. By [ENO05, Proposition 8.27] we have that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer. By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$. The second item follows since for a modular category we have $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$, because the group associated to ${\mathcal{C}_{\mathrm{pt}}}$ is the universal grading group by [EGNO15, Lemma 8.22.5]. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$. On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})\big{)}$.
Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$, and by [EGNO15, Proposition 8.14.6], $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $i\in\mathcal{I}$. ∎
@@
Since $\gcd(d_{i}^{2},d)=1$ and $d$ is square-free, it follows from Equation (1.13) that $d$ divides $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Now, consider a pointed fusion subcategory ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ such that $\mathrm{FPdim}({\mathcal{D}})=d$. The existence of such a subcategory is straightforward to demonstrate because the associated groups are abelian, given that the category is braided. According to [EGNO15, Theorem 8.21.5], with $\mathcal{C}^{\prime}=\mathrm{Vec}$, we have $\mathrm{FPdim}({\mathcal{D}}^{\prime})=m$, but $\gcd(d,m)=1$. Consequently, by Lagrange’s theorem [EGNO15, Theorem 7.17.6], ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$, which implies that ${\mathcal{D}}$ is non-degenerate, as stated in [EGNO15, Corollary 8.20.10]. Thus, by [EGNO15, Theorem 8.21.4], $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, and ${\mathcal{D}}^{\prime}$ is also non-degenerate. ∎
@@
Note that Theorem 1.21 improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
@@
Immediate from Theorem 1.21, we just need to check that $gcd(d_{i}^{2},d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a factor of the square-free part. ∎
Corollary 10.9.
An integral modular fusion category $\mathcal{C}$ can always be decomposed into ${\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, where both ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$ are modular, ${\mathcal{D}}$ is pointed, and $\mathrm{FPdim}({\mathcal{D}})$ is the maximal square-free part of $\mathrm{FPdim}(\mathcal{C})$.
This means that the classification of integral modular fusion categories $\mathcal{C}$ simplifies to those where $\mathrm{FPdim}(\mathcal{C})$ lacks a square-free part, up to a Deligne tensor product with a pointed modular fusion category.
Corollary 10.10.
For an integral perfect modular fusion category $\mathcal{C}$, $\mathrm{FPdim}(\mathcal{C})$ has no square-free part. In other words, if a prime $p$ divides $\mathrm{FPdim}(\mathcal{C})$, then $p^{2}$ also divides $\mathrm{FPdim}(\mathcal{C})$.
Corollary 10.10 obviously extends to the weakly-integral case due to the following:
@@
Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, because $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8], see also Remark 11.13. Moreover, [EGNO15, Exercise 9.6.12] asserts that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
@@
If $\mathcal{C}$ is a perfect, integral, even-dimensional modular fusion category, then 444 divides $\mathrm{FPdim}(\mathcal{C})$.
10.2. On Statements (1.25) and (1.26) and Conjecture (1.24)
In [LPR23, Theorem 5.6] it was shown that Statements (1.25) and (1.26) are equivalent. We will show that any of these statements implies Conjecture 1.24 formulated in [CP22].
@@
The existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
@@
Let $\mathcal{C}$ be a perfect modular fusion category with an odd dimension. By Proposition 10.11, $\mathcal{C}$ must be integral.
[…]
However, due to the finite number of isomorphism classes of simple objects, $\mathcal{C}$ must contain at least one minimal subcategory. Let us call this minimal subcategory ${\mathcal{D}}$. If ${\mathcal{D}}$ is itself a modular fusion category, then ${\mathcal{D}}$ serves as an example of a non-pointed, simple, integral modular fusion category.
If ${\mathcal{D}}$ is not modular, then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is non-trivial according to [EGNO15, Corollary 8.20.10], yet it is symmetric by design. But as discussed earlier, such symmetric subcategories cannot exist, leading to a contradiction. ∎